The solution of equation log7 log5 (√x + 5) + √x) = 0 is . . . . . . . . . . . . . . . .

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. log7 log5 (√x + 5) + √x) = 0
⇒ log5 (√x + 5) + √x) = 1 NOTE THIS STEP
⇒ √x + 5 + √x = 5 ⇒ x + 5 = 25 + x - 10√x
⇒ 2 = √x ⇒ x = 4 which satisfies the given equation.